OpenStudy (anonymous):

Jimmy's Chocolate shop sells two different fudge bars. The dark chocolate fudge bars sell for $5 per pound and the milk chocolate fudge bars sell for $4 per pound. You purchased a combination of these bars for your family and friends and your bill came to $150. This scenario can be modeled by the equation 5d + 4m = 150. Which of the following is not a possible combination of dark chocolate fudge bars and milk chocolate fudge bars? A) 6 dark bars and 30 milk bars B) 10 dark bars and 25 milk bars C) 12 dark bars and 23 milk bars D) 14 dark bars and 20 milk bars

OpenStudy (swissgirl):

Well our equation is 5d+4m=150 where d=dark chocolate and m=milk chocolate

OpenStudy (swissgirl):

Ok so they gave us a list of options and we have to plug them into the equation, The option that doesnt equal 150 is the answer since that isnt a possible combination 5(6)+4(30)=30+120=150 so option#1 equals 150 so its a possible combination

OpenStudy (anonymous):

so the answer is a? @swissgirl

OpenStudy (swissgirl):

5(10)+4(25)=50+100=150 Option#2 equals 150 so its a possible combination

OpenStudy (swissgirl):

nope its not a since it equals 150 We have to find the option that doesnt equal 150

OpenStudy (anonymous):

ooooh okayy @swissgirl

OpenStudy (swissgirl):

5(12)+4(23)=60+92 = 152 The third option doesnt equal 150 so this is our answer